The highest recorded temperature on Krypton was 163[sup]◦[/sup]F.&
The lowest recorded temperature was −134[sup]◦[/sup]F.&
Express the low temperature extreme indegrees Celsius.
Answer in units of& [sup]◦[/sup]C.
1
Expert's answer
2011-08-13T08:14:02-0400
Common formula of conversion:
C = 5/9 (F-32)

C1 = 5/9 (163 - 32) = 72.78 ≈ 73C
C2 = 5/9 (-134 - 32) = 92.22 ≈ 92C